# The disk precession in a Be star-magnetar binary and its application to the rotation measure of FRB 20201124A

### v1

@@ 3.2 Disk precession
The Be star disk precesses around a precession axis, which can be defined by a unit vector $\vec{e}_{\rm{pre}}$ and forms an angle $\varphi$ with the $z$ axis. To simplify the scenario, $\varphi$ is constrained in the $x$-$z$ plane. $\theta$ is the angle between $\vec{e}_{\rm{disk}}$ and $\vec{e}_{\rm{pre}}$. If the period of precession is $P_{\rm{pre}}$, $\vec{e}_{\rm{pre}}$ and $\vec{e}_{\rm{disk}}$ can be expressed as follows:
@@
where the $\lambda_{1}$ or $\lambda_{2}$ are the angles between $\vec{e}_{\rm{LoS}}$ and $\vec{e}_{x}$ or $\vec{e}_{z}$.
4 Application to FRB 20201124A
In this section, we will use the precession disk model of the B star-magnetar binary in Section 3.2 to explain the variations in the RM data of FRB 20201124A.
4.1 Model parameters
In our model described in Section 3, the parameters are as follows
@@ 4.1 Model parameters
For the Stellar parameters, their values are fixed in our calculations. For a B type star, its mass range is poorly constrained, spanning from a handful to a dozen of $M_{\odot}$. In this study, we fix $M\star$ as 8 $M_{\odot}$. Since the radius of B type stars ranges from a few to ten $R_{\odot}$, we adopt $R_{\star}=5\,R_{\odot}$. For the magnetar mass $M_{\rm{NS}}$, we use a typical NS mass of 1.6 $M_{\odot}$.
@@
For the Disk properties, the isothermal temperature, the mean molecular weight, and the initial magnetic field in the disk are assumed to be $T=20,000\,\rm{K}$, $\mu=1$, and $\rm{B_{0}}=10\,\rm{G}$, respectively, referring to Wang_2022NatCo..13.4382W. Since the fitting is sensitive to the density profile of the disk, we set $\rho_{0}$ and $\beta$ as free parameters to achieve a better fitting result.
@@
For the orbital parameters, i.e., $P_{\rm{orb}}$ and $\varepsilon$, we fix $P_{\rm{orb}}$ but treat $\varepsilon$ as a free parameter. According to the RM data of FRB 20201124A and our model, we anticipate that the maximum dips of E1 ($\sim$ MJD 59337) and E3 ($\sim$ MJD 59627) as seen in Figure 2 are very likely in the same phase of the magnetar orbital motion. Hence, our model expects that the time span between the two dips is an integer multiple of the orbital period $P_{\rm{orb}}$ and estimates the period to be $P_{\rm{orb}}=73$ days. With $P_{\rm{orb}}=73$ days, 4 orbital periods have been covered by the current FAST observations (see Figure 2). Since the orbital eccentricity $\varepsilon$ can significantly affect the profile of the RM variations, it is treated as a free parameter.
@@
For the parameters of disk precession and LoS, since they are directly related to precession and RM integration, all of them are set as free parameters. In conclusion, our fitting parameters are $\varepsilon$, $\rho_{0}$, $\beta$, $\varphi$, $\lambda_{1}$, $\lambda_{2}$, $\theta$, and $\omega$.
4.2 Calculation methods and results
@@
To compute the LoS integral of RM, we apply the method developed by Wang_2022NatCo..13.4382W: At each orbital phase, we take 1000 points along the LoS starting from the magnetar’s position, with a distance of $1\,R_{\odot}$ between two adjacent points. For each calculation point, the local plasma density and the magnetic field are calculated from Equations 1 and 4, respectively, and the expected $\Delta\rm{RM}$ is calculated from Equation 6.
@@
Figure 2: The RM variation of FRB 20201124A in three FAST observation episodes, i.e., $\Delta\rm{RM}_{E1}$ (blue dots), $\Delta\rm{RM}_{E2}$ (maroon dots) and $\Delta\rm{RM}_{E3}$ (dark green dots). The ranges of observed RM of E1, E2 and E3 are shown as the blue, maroon, and green shadow regions, respectively. The fitting of the model presented in Section 3 is shown as the red curve, while the parameters of the fitting is shown in Table 1.
@@
The orbital eccentricity $\varepsilon$ mainly affects the ratio of the peak and dip durations, as well as the amplitude of the $\Delta\rm{RM}$ curve. The larger $\varepsilon$, the greater the proportion of the peak in one period, and the more drastic the variation in the $\Delta\rm{RM}$ curve. The base mass density $\rho_{0}$ and the density radial profile $\beta$ can also influence the variation amplitude of the $\Delta\rm{RM}$ curve: increasing with increasing $\rho_{0}$ or decreasing $\beta$. For $\theta$, larger values lead to an increase in both the duration of the dip and its amplitude. $\varphi$ variations induce changes in the shape of the $\Delta\rm{RM}$ curve; positive $\varphi$ leads to a taller and wider peak and a tiny dip, while a more negative $\varphi$ results in a sharper dip (positive and negative $\varphi$ correspond to the positive and negative projections of $\vec{e}_{\rm{pre}}$ onto the x-axis in Figure 1, respectively). For $\omega$, it regulates how quickly the overall amplitude of the $\Delta\rm{RM}$ curve oscillates over each orbital period. The larger $\omega$, the faster the oscillation. For the LoS, variations in $\lambda_{1}$ significantly affect the proportion and amplitude of peaks and dips within one orbital period. A smaller $\lambda_{1}$ can result in a wider and deeper dip as well as a narrower and lower peak. For $\lambda_{2}$, its small perturbations can lead to significant changes in the amplitude of the $\Delta\rm{RM}$ curve. As $\lambda_{2}$ decreases, the amplitude of the $\Delta\rm{RM}$ curve increases — with the peak part increasing far more significantly than the dip part within an orbital period.
@@
We fit the $\Delta\rm{RM}_{\rm{E1}}$, $\Delta\rm{RM}_{\rm{E2}}$ and $\Delta\rm{RM}_{\rm{E3}}$ data with our model. We obtained a set of parameters that are reasonable and well fit the RM variation data: $\varepsilon=0.69$, $\rho_{0}=3.49\times 10^{-15}\,\rm{g\,cm^{-3}}$, $\beta=4.09$, $\theta=0.07\,\rm{rad}$ ($\sim 4^{\circ}$), $\varphi=-0.7\,\rm{rad}$, $\omega=8.01\times 10^{-3}\,\rm{rad\,day^{-1}}$ (corresponding to a precession period of 784.71 days), $\lambda_{1}=-0.55\,\rm{rad}$, and $\lambda_{2}-\pi=-9.64\times 10^{-2}\,\rm{rad}$ (since $\lambda_{2}$ is very close to $\pi$, we adopt $\lambda_{2}-\pi$ to represent its magnitude). The results are presented in Table 1, and the fitting curve plotted using the best-fit parameters is shown in Figure 1 as the red curve.
@@
5 Discussions and Conclusions
5.1 Discussions
@@
Unlike FRB 20201124A, these FRB sources do not have enough data from observations, making the constraint on disk precession loose. We apply our model to three of these FRBs, i.e. FRB 20180301A, FRB 20190520B, and FRB 20191106C, and also offer a prediction for their missing part of RM observations. The fittings and predictions of these sources are shown in Figure 3.
@@
Figure 3: Three FRB sources with notable RM variation, each with model fitting RM curves of $P_{\rm{prec}}=10\times P_{\rm{orb}}$ (blue curve) and without precession (dotted red curve). The two curves in each subplot share a same set of parameters except $\theta$ and $\omega$, since these two parameters are both 0 for the no-precession case. Evolution of this two curves are shown for five orbital periods in each subplot. The observed $\Delta\rm{RM}$s is shown as the gray dots. The parameters of each FRB source are shown in Table 2.
In this work, we assume an ideal uniform precession of the Be star with a rigid decretion disk. The case would be, However, more complex in practice. In fact, whether a disk can behave as a rigid body depends on whether the interactions between its different parts are sufficiently strong — such as bending waves, viscous stresses, and self-gravity (Papaloizou_&Pringle_1983MNRAS.202.1181P; Papaloizou_&_Lin_1995ApJ...438..841P; Papaloizou_&_Terquem_1995MNRAS.274..987P; Ogilvie_1999MNRAS.304..557O; Ogilvie_2006MNRAS.365..977O; Bate_2000MNRAS.317..773B; Lubow_&_Ogilvie_2000ApJ...538..326L; Batygin_2012Natur.491..418B; Tremaine_2014MNRAS.441.1408T). Furthermore, according to Lai_2014MNRAS.440.3532L, the circumstellar disk of a massive star in a binary system might undergo precessional evolution, which is driven by the gravitational torques exerted by the companion star tidally. Larger stellar obliquity, i.e., the misalignment between the equatorial plane of the star and the orbital plane, might be induced by such evolution of the disk precession. However, such disk evolution is expected to be significant over a timescale of mega-years, so we do not consider it in our model.
According to the calculation of a rigid disk precession induced by a companion star in Lai_2014MNRAS.440.3532L, we estimate the precession period using the binary and disk parameters in Section 4.2. The disk with the density profile in Equation 1 can be regarded as a thin disk that spans from the inner radius $R_{\rm{in}}=R_{\star}$ to $R_{\rm{out}}=\eta R_{\star}$ with surface density:
@@
$M_{d}=\int dM_{d}=2\pi\int_{R_{\rm{in}}}^{R_{\rm{out}}}r\Sigma(r)dr=\frac{4\pi\Sigma_{0}R_{\star}^{4}}{7-2\beta}(\eta^{7/2}-1),$ (15)
and for the disk angular momentum vector $\vec{L_{d}}=L_{d}\,\vec{e_{L_{d}}}$ (where $\vec{e_{L_{d}}}$ is the unit vector), can be calculated as
$L_{d}=\int rM_{d}(r)v_{k}(r)dr=\frac{2\pi\Sigma_{0}R_{\star}^{5/2}\sqrt{GM_{\star}}}{4-\beta}(\eta^{4-\beta}-1).$ (16)
@@
$\vec{T_{d}}$ $=-\int^{R_{\rm{out}}}_{R_{\rm{in}}}dM_{d}\frac{3GM_{\rm{NS}}}{4a_{b}^{3}}\cos\theta\,\vec{e_{L_{b}}}\times\vec{e_{L_{d}}}$ (17)
$=\frac{3\pi GM_{\rm{NS}}\Sigma_{0}R_{\star}^{4}\cos\theta}{(11-2\beta)a_{b}^{3}}(\eta^{11/2-\beta}-1)\,\vec{e_{L_{b}}}\times\vec{e_{L_{d}}}$
$=\omega\vec{e_{L_{b}}}\times\vec{L_{d}},$
@@
where $\vec{e_{L_{b}}}$ is the unit vector along the binary angular momentum axis, $a_{b}$ is the binary separation, $\theta$ is the inclination angle between $\vec{e_{L_{b}}}$ and $\vec{e_{L_{d}}}$, and $\omega$ is the angular velocity of the disk precession. Thus, the precession period of the disk, $P_{\rm{prec}}$, can be calculated as
$P_{\rm{prec}}$ $=2\pi/\omega=2\pi/|\vec{T_{d}}/\vec{L_{d}}|$ (18)
$=\frac{4\pi{a_{b}}^{3}}{3M_{\rm{NS}}R_{\star}^{3/2}cos\theta}\sqrt{\frac{M_{\star}}{G}}(\frac{11-2\beta}{4-\beta})(\frac{\eta^{4-\beta}-1}{\eta^{11/2-\beta}-1}).$
According to Section 4.2, $\beta=4$ is reasonable for the case of FRB 20201124A. When $\beta=4$, one has
$P_{\rm{prec}}^{\prime}$ $\simeq\frac{4\pi a_{b}^{3}}{M_{\rm{NS}}R_{\star}^{3/2}cos\theta}\sqrt{\frac{M_{\star}}{G}}\frac{ln(\eta)}{\eta^{3/2}-1}.$ (19)
@@
$\eta$ shows an inverse correlation with the precession period $P_{\rm{prec}}^{\prime}$. We adopt the parameter values from Table 1, i.e., $\theta=0.07\,\rm{rad}$, $M_{\rm{NS}}=1.6M_{\odot}$, $M_{\star}=8M_{\odot}$, and $R_{\star}=5R_{\odot}$. We take the time averaged separation for $a_{b}$:
$\langle a_{b}\rangle_{t}\ =\frac{1}{P_{\rm{orb}}}\int_{0}^{P_{\rm{orb}}}a_{b}(t)dt=a(1+e^{2}/2),$ (20)
@@
where $a$ and $P_{\rm{orb}}$ are the semi-major axes of the binary orbit. Here, we adopt $P_{\rm{orb}}=73\,\rm{day}$ and $e=0.69$ from Table 1 to calculate $a$ (see Equation 5) and $\langle a_{b}\rangle_{t}\$.
@@
From the above formulae, we calculate that $\eta$ spans the range $[100,200]$, corresponding to $P_{\rm{prec}}^{\prime}$ in the range $[498,1220]$ days. The $\sim 900$-day period in Section 3 corresponds to $\eta\sim 141$. To estimate the impact of $\beta$ on the precession period, we fix $\eta$ to 125 in Equation 18. A $P_{\rm{prec}}$ in the range $[500,1200]$ days, corresponding to $\beta$ in the range $[3.71,4.21]$, and in this calculation $\beta=4$ is not included, as $\beta=4$ is a singularity in Equation 18 (an increase in $\beta$ leads to a corresponding increase in $P_{\rm{prec}}$). This is comparable with the $\beta$ in Table 1.
@@
This means that using the model in Section 3 with reasonable $\eta$ and $\beta$, one can adequately explain the expected $\sim 785$-day precession period in Section 4.2.
5.2 Conclusions

### v2

@@ III.2 Disk precession
The Be star disk precesses around a precession axis, which can be defined by a unit vector $\vec{e}_{\rm{pre}}$ and forms an angle $\varphi$ with the $z$ axis. To simplify the scenario, $\varphi$ is constrained to the $x$-$z$ plane. $\theta$ is the angle between $\vec{e}_{\rm{disk}}$ and $\vec{e}_{\rm{pre}}$. If the period of precession is $P_{\rm{pre}}$, $\vec{e}_{\rm{pre}}$ and $\vec{e}_{\rm{disk}}$ can be expressed as follows:
@@
where $\lambda_{1}$ or $\lambda_{2}$ are the angles between $\vec{e}_{\rm{LoS}}$ and $\vec{e}_{x}$ or $\vec{e}_{z}$.
IV Application to FRB 20201124A
In this section, we will use the precession disk model of the B star-magnetar binary in Section III.2 to explain the variations in FRB 20201124A RM data.
IV.1 Model parameters
In our model described in Section III, the parameters are as follows
@@ IV.1 Model parameters
For stellar parameters, their values are fixed in our calculations. For a B type star, its mass range is poorly constrained, spanning from a handful to a dozen $M_{\odot}$. In this study, we fix $M\star$ as 8 $M_{\odot}$. Since the radius of B-type stars ranges from a few to ten $R_{\odot}$, we adopt $R_{\star}=5\,R_{\odot}$. For the magnetar mass $M_{\rm{NS}}$, we use a typical NS mass of 1.6 $M_{\odot}$.
@@
For disk properties, the isothermal temperature, the mean molecular weight, and the initial magnetic field in the disk are assumed to be $T=20,000\,\rm{K}$, $\mu=1$, and $\rm{B_{0}}=10\,\rm{G}$, respectively, referring to Wang et al. (2022). Since the fitting is sensitive to the density profile of the disk, we set $\rho_{0}$ and $\beta$ as free parameters to achieve a better fitting result.
@@
For orbital parameters, i.e., $P_{\rm{orb}}$ and $\varepsilon$, we fix $P_{\rm{orb}}$ but treat $\varepsilon$ as a free parameter. According to the RM data of FRB 20201124A and our model, we anticipate that the maximum dips of E1 ($\sim$ MJD 59337) and E3 ($\sim$ MJD 59627) as seen in Figure 2 are very likely in the same phase of magnetar orbital motion. Hence, our model expects that the time span between the two dips is an integer multiple of the orbital period $P_{\rm{orb}}$ and estimates the period to be $P_{\rm{orb}}=73$ days. With $P_{\rm{orb}}=73$ days, four orbital periods have been covered by the current FAST observations (see Figure 2). Since the orbital eccentricity $\varepsilon$ can significantly affect the profile of the RM variations, it is treated as a free parameter.
@@
For parameters of disk precession and LoS, as they are directly related to precession and RM integration, all of them are set as free parameters. In conclusion, our fitting parameters are $\varepsilon$, $\rho_{0}$, $\beta$, $\varphi$, $\lambda_{1}$, $\lambda_{2}$, $\theta$, and $\omega$.
IV.2 Results
@@
To compute the LoS integral of RM, we apply the method developed by Wang et al. (2022): At each orbital phase, we take 1000 sample points along the LoS starting from the magnetar’s position, with a distance of $1\,R_{\odot}$ between two adjacent points. At each calculation point, the local plasma density and the magnetic field are calculated from Equations (1) and (4), respectively, and the expected $\Delta\rm{RM}$ is calculated from Equation (6).
@@
Figure 2: The RM variations of FRB 20201124A in three FAST observation episodes, i.e., $\Delta\rm{RM}_{E1}$ (blue dots), $\Delta\rm{RM}_{E2}$ (maroon dots) and $\Delta\rm{RM}_{E3}$ (dark green dots). The observed RM ranges of E1, E2 and E3 are shown as the blue, maroon, and green shadow regions, respectively. The fit of the model presented in Section III is shown as the red curve, and the corresponding fitting parameters are listed
in Table 1.
@@
The orbital eccentricity $\varepsilon$ mainly affects the ratio between the peak and dip durations, as well as the amplitude of the $\Delta\rm{RM}$ curve. A larger $\varepsilon$ increases the relative duration of the peak within one orbital period and leads to more pronounced variations in the $\Delta\rm{RM}$ curve. The base mass density $\rho_{0}$ and the density radial profile $\beta$ can also influence the variation amplitude of the $\Delta\rm{RM}$ curve: increasing with increasing $\rho_{0}$ or decreasing $\beta$. For $\theta$, larger values lead to an increase in both the duration of the dip and its amplitude. $\varphi$ variations induce changes in the shape of the $\Delta\rm{RM}$ curve; positive $\varphi$ leads to a taller and wider peak and a tiny dip, while a more negative $\varphi$ results in a sharper dip (positive and negative $\varphi$ correspond to the positive and negative projections of $\vec{e}_{\rm{pre}}$ onto the x-axis in Figure 1, respectively). For $\omega$, it regulates how quickly overall amplitude of the $\Delta\rm{RM}$ curve oscillates over each orbital period. The larger $\omega$, the faster the oscillation. For the LoS, variations in $\lambda_{1}$ significantly affect proportion and amplitude of peaks and dips within one orbital period. A smaller $\lambda_{1}$ can result in a wider and deeper dip as well as a narrower and lower peak. For $\lambda_{2}$, its small perturbations can lead to significant changes in amplitude of the $\Delta\rm{RM}$ curve. As $\lambda_{2}$ decreases, the amplitude of the $\Delta\rm{RM}$ curve increases — with the peak part increasing far more significantly than the dip part within an orbital period.
@@
We fit $\Delta\rm{RM}_{\rm{E1}}$, $\Delta\rm{RM}_{\rm{E2}}$ and $\Delta\rm{RM}_{\rm{E3}}$ data with our model. We obtained a set of parameters that are reasonable and well fit the RM variation data: $\varepsilon=0.69$, $\rho_{0}=3.49\times 10^{-15}\,\rm{g\,cm^{-3}}$, $\beta=4.09$, $\theta=0.07\,\rm{rad}$ ($\sim 4^{\circ}$), $\varphi=-0.7\,\rm{rad}$, $\omega=8.01\times 10^{-3}\,\rm{rad\,day^{-1}}$ (corresponding to a precession period of 784.71 days), $\lambda_{1}=-0.55\,\rm{rad}$, and $\lambda_{2}-\pi=-9.64\times 10^{-2}\,\rm{rad}$ (since $\lambda_{2}$ is very close to $\pi$, we adopt $\lambda_{2}-\pi$ to represent its magnitude). The results are presented in Table 1, and the fitting curve plotted using the best-fit parameters is shown in Figure 1 as the red curve.
@@
V Discussions and Conclusions
V.1 Discussions
V.1.1 Other FRBs
@@
Unlike FRB 20201124A, these FRB sources do not yet have sufficient observational data, making constraint on disk precession loose. We apply our model to three of these FRBs, i.e. FRB 20180301A, FRB 20190520B, and FRB 20191106C, and also offer a prediction for their missing part of RM observations. The $\Delta\rm{RM}$ data of FRB 20180301A, FRB 20190520B, and FRB 20191106C are obtained from Kumar et al. (2023), Anna-Thomas et al. (2023), and Ng et al. (2025), respectively. The fittings and predictions of these sources are shown in Figure 3.
@@
Figure 3: Three FRB sources with notable RM variation, each with model fitting RM curves of $P_{\rm{prec}}=10\times P_{\rm{orb}}$ (blue curve) and without precession (dotted red curve). The two curves in each subplot share the same set of parameters except $\theta$ and $\omega$, since these two parameters are both 0 for the no-precession case. Evolutions of these two curves are shown for five orbital periods in each subplot. The observed $\Delta\rm{RM}$s are shown as the gray dots. The parameters of each FRB source are shown in Table 2.
V.1.2 Dynamical Analysis of Disk Precession
In this work, we assume an ideal uniform precession of the Be star with a rigid decretion disk. However, the case would be more complex in practice. In fact, whether a disk can behave as a rigid body depends on whether the interactions between its different parts are sufficiently strong — such as bending waves, viscous stresses, and self-gravity (Papaloizou and Terquem, 1995; Ogilvie, 1999, 2006; Bate et al., 2000; Batygin, 2012; Tremaine and Davis, 2014). Furthermore, according to Lai (2014), the circumstellar disk of a massive star in a binary system might undergo precessional evolution, which is driven by the gravitational torques exerted by the companion star tidally. Larger stellar obliquity, i.e., the misalignment between the equatorial plane of the star and the orbital plane, might be induced by such evolution of the disk precession. However, such disk evolution is expected to be significant over a timescale of mega-years, so we do not consider it in our model.
Following the calculation of a rigid disk precession induced by a companion star in Lai (2014), we estimate the precession period using the binary and disk parameters in Section IV.2. The disk with the density profile in Equation (1) can be regarded as a thin disk that spans from the inner radius $R_{\rm{in}}=R_{\star}$ to $R_{\rm{out}}=\eta R_{\star}$ with surface density:
@@ V.1.2 Dynamical Analysis of Disk Precession
$M_{\rm d}=\int dM_{\rm d}=2\pi\int_{R_{\rm{in}}}^{R_{\rm{out}}}r\Sigma(r)dr=\frac{4\pi\Sigma_{0}R_{\star}^{4}}{7-2\beta}(\eta^{7/2}-1),$ (15)
and for the disk angular momentum vector $\vec{L}_{\rm d}=L_{\rm d}\,\vec{e}_{{L}_{\rm d}}$ (where $\vec{e}_{L_{\rm d}}$ is the unit vector) can be calculated as
$L_{\rm d}=\int rM_{\rm d}(r)v_{k}(r)dr=\frac{2\pi\Sigma_{0}R_{\star}^{5/2}\sqrt{GM_{\star}}}{4-\beta}(\eta^{4-\beta}-1).$ (16)
@@
$\vec{T}_{\rm d}$ $=-\int^{R_{\rm{out}}}_{R_{\rm{in}}}dM_{\rm d}\frac{3GM_{\rm{NS}}}{4a_{\rm b}^{3}}\cos\theta\,\vec{e}_{L_{\rm b}}\times\vec{e}_{L_{\rm d}}$ (17)
$=\frac{3\pi GM_{\rm{NS}}\Sigma_{0}R_{\star}^{4}\cos\theta}{(11-2\beta)a_{\rm b}^{3}}(\eta^{11/2-\beta}-1)\,\vec{e}_{L_{\rm b}}\times\vec{e}_{L_{\rm d}}$
$=\omega\vec{e}_{L_{\rm b}}\times\vec{L}_{\rm d},$
@@
where $\vec{e}_{L_{\rm b}}$ is the unit vector along the binary angular momentum axis, $a_{\rm b}$ is the binary separation, $\theta$ is the inclination angle between $\vec{e}_{L_{\rm b}}$ and $\vec{e}_{L_{\rm d}}$, and $\omega$ is the angular velocity of the disk precession. Thus, the precession period of the disk, $P_{\rm{prec}}$, can be calculated as
$P_{\rm{prec}}$ $=2\pi/\omega=2\pi/|\vec{T}_{\rm d}/\vec{L}_{\rm d}|$ (18)
$=\frac{4\pi{a_{\rm b}}^{3}}{3M_{\rm{NS}}R_{\star}^{3/2}\rm{cos}\theta}\sqrt{\frac{M_{\star}}{G}}(\frac{11-2\beta}{4-\beta})(\frac{\eta^{4-\beta}-1}{\eta^{11/2-\beta}-1}).$
According to Section IV.2, $\beta=4$ is reasonable for the case of FRB 20201124A. When $\beta=4$, one has
$P_{\rm{prec}}$ $\simeq\frac{4\pi a_{\rm b}^{3}}{M_{\rm{NS}}R_{\star}^{3/2}\rm{cos}\theta}\sqrt{\frac{M_{\star}}{G}}\frac{\ln(\eta)}{\eta^{3/2}-1}.$ (19)
@@
$\eta$ shows an inverse correlation with the precession period $P_{\rm{prec}}$. We adopt the parameter values from Table 1, i.e., $\theta=0.07\,\rm{rad}$, $M_{\rm{NS}}=1.6M_{\odot}$, $M_{\star}=8M_{\odot}$, and $R_{\star}=5R_{\odot}$. We take the time averaged separation for $a_{b}$:
$\langle a_{\rm b}\rangle_{t}\ =\frac{1}{P_{\rm{orb}}}\int_{0}^{P_{\rm{orb}}}a_{\rm b}(t)dt=a(1+e^{2}/2),$ (20)
@@
where $a$ and $P_{\rm{orb}}$ are the semi-major axes of the binary orbit. Here, we adopt $P_{\rm{orb}}=73\,\rm{day}$ and $e=0.69$ from Table 1 to calculate $a$ (see Equation (5)) and $\langle a_{b}\rangle_{t}\$.
@@
From the above formulae, we calculate that $\eta$ spans the range $[100,200]$, corresponding to $P_{\rm{prec}}$ in the range $[498,1220]$ days. The $\sim 800$-day period in Section III corresponds to $\eta\sim 141$. To estimate the impact of $\beta$ on the precession period, we fix $\eta$ to 125 in Equation (18). A $P_{\rm{prec}}$ in the range $[500,1200]$ days, corresponding to $\beta$ in the range $[3.71,4.21]$, and in this calculation $\beta=4$ is not included, as $\beta=4$ is a singularity in Equation (18) (an increase in $\beta$ leads to a corresponding increase in $P_{\rm{prec}}$). This is comparable to $\beta$ in Table 1.
@@
This indicates that, with reasonable values of $\eta$ and $\beta$, the model in Section III can adequately account for the expected $\sim 785$-day precession period in Section IV.2.
V.1.3 Absorption of FRBs by the Be Star Disk
When an FRB penetrates the decretion disk of a Be star, the absorption process by the disk material may occur (Wang et al., 2022; Zhang, 2023). To make sure that the FRB can reach the observer, it is essential to discuss the absorption process.
The free-free absorption can be important when the medium density is high along the propagation path of the FRB. The optical depth of the free-free absorption at frequency $\nu$, i.e. $\tau^{\rm{ff}}_{\nu}$, can be calculated as follows:
$\tau^{\rm{ff}}_{\nu}$ $=\int_{\rm{LoS}}\alpha_{\nu}^{\rm{ff}}\,dl$ (21)
[…]
where $\alpha_{\nu}^{\rm{ff}}$ is the free-free absorption coefficient, and $\nu$ is the frequency of FRB signal. Here, $T$ and $\nu$ are in units of Kelvin and GHz, respectively. Using the lowest observation frequency of 1 GHz and $T=20,000K$ from Section IV.1, with parameters of the model from Table 1, the calculated $\tau^{\rm{ff}}_{\nu}$ is below $5\times 10^{-4}$ (see Figure 4), which is much less than 1. Thus, in the case of FRB 20201124A, the disk has no effect on the detectability of the radio signal.
Figure 4: The optical depth of free-free absorption contributed by the Be star decretion disk in the case of FRB 20201124A. The evolution time of $\tau^{\rm{ff}}_{\nu}$ consistent with that in Figure 2, showing the variation of $\tau^{\rm{ff}}_{\nu}$ over 5 orbital periods.
For the other three FRBs in Section V.1.1, we also calculate each of their $\tau^{\rm{ff}}_{\nu}$ using the parameters in Table 2. The observation frequency ranges of FRB 20180301A, FRB 20180520B, and FRB 20191106C are $\sim$ 0.7 to 1.5 GHz, $\sim$ 0.7 to 8 GHz, and $\sim$ 0.4 to 0.8 GHz, respectively (Kumar et al., 2023; Anna-Thomas et al., 2023; Ng et al., 2025). Since $\tau^{\rm{ff}}_{\nu}\propto\nu^{-2.1}$, we adopt 0.7 GHz in the calculation of FRB 20180301A and FRB 20180520B, and 0.4 GHz for FRB 20191106C. It turns out that their free-free absorption optical depths are far less than unity.
The synchrotron absorption may also play a role. However, given that our work assumes scenarios after the FRB has been emitted from the emitter environment, we do not consider absorption effects associated with the emitter environment here. On the other hand, since we adopt a cold disk (20,000 K) in our work that lacks relativistic electrons, we can neglect the synchrotron absorption contributed by the disk as the FRB propagates through it.
V.1.4 The magnetic field in the stellar outflow
For an early-type massive star, the magnetic field is dipolar within a certain region that extends from the stellar surface (Usov and Melrose, 1992). In this region, the stellar outflow moves along the dipolar field and is confined by these closed magnetic field lines near the equatorial plane. This region is defined by the Alfvénic surface, where the ram pressure of the outflow is equal to the characteristic value of the magnetic stress components (Mestel and Spruit, 1987; Usov and Melrose, 1992):
$\frac{1}{2}\rho(R_{\rm A})v_{\rm{w}}^{2}=\frac{\mathrm{B}_{\rm ms}(R_{\rm A})^{2}}{8\pi},$ (22)
where $\rho(r)=\rho_{0}(r/R_{\star})^{-\beta}$ is the density profile of the outflow disk in the equatorial plane according to Equation (1), $R_{\rm A}$ is the radius of the Alfvénic surface (i.e., the Alfvén radius), and $v_{\rm{w}}$ is the wind velocity. The wind velocity $v_{\rm{w}}$ can be estimated as the terminal speed of the wind $v_{\rm{w,\infty}}$. $v_{\rm w,\infty}$ of $\sim 10^{7}-10^{8}$ cm/s is typical for B-type stars from observations (Snow, 1981; Krtička, 2014). $\mathrm{B}_{\rm ms}(r)$ denotes the magnetic field strength of the massive star in the out-flowing gas (Usov and Melrose, 1992; Melatos et al., 1995; Zhao et al., 2023):
$\mathrm{B}_{\rm ms}(r)=\begin{cases}\mathrm{B}_{\rm s}(\frac{r}{R_{\star}})^{-3},&(\mathrm{dipolar},\,R_{\star}<r<R_{\rm A})\\ \,\\ \mathrm{B}_{\rm s}(\frac{R_{\rm A}}{R_{\star}})^{-3}(\frac{r}{R_{\rm A}})^{-2},&(\mathrm{radial},\,R_{\rm A}<r<\frac{v_{\rm{w,\infty}}}{v_{\rm{rot}}}R_{\star})\\ \,\\ \mathrm{B}_{\rm s}(\frac{R_{\rm A}}{R_{\star}})^{-3}(\frac{r}{R_{\rm A}})^{-1},&(\mathrm{toroidal},\,r>\frac{v_{\rm{w,\infty}}}{v_{\rm{rot}}}R_{\star})\end{cases},$ (23)
where $\mathrm{B}_{\rm s}$ is the magnetic field strength at the stellar surface, and $v_{\rm{rot}}$ denotes the surface rotation velocity of Be star. For typical values of surface rotation velocity for O or B-type stars are $v_{\rm{rot}}\sim 0.15\,v_{\rm{w,\infty}}$. As indicated in Equation (23), the magnetic field beyond the Alfvén surface becomes radial due to the drag effect of the stellar wind. Eventually, the magnetic field changes from radial to toroidal at a larger distance, which is consistent with Equation (4). Thus, $R_{\rm A}$ can be obtained from Equations (22) and (23).
For FRB 20201124A, taking into account that the toroidal magnetic field in Equation (4) and 23 is the same, we calculate a surface magnetic field strength $\mathrm{B}_{\rm s}=40\,\rm{G}$, with Alfvén radius $R_{\rm A}\sim 2R_{\star}\,(10\,R_{\odot})$ ($v_{\rm{w}}=10^{8}$ cm/s is adopted in calculation). In comparison, using the LoS and orbital parameters of FRB 20201124A from Table 1, the perpendicular distance between the FRB’s LoS and the Be star, $D_{\perp}$, is calculated to be about $45\sim 250\,R_{\odot}$, which is much larger than $R_{\rm A}$, and also exceeds the range dominated by the radial magnetic field (the calculation results can be seen in Table 3. Thus, the dipolar magnetic field of the Be star does not contribute to the RM of FRB 20201124A, and the magnetic field along the LoS of FRB 20201124A can be regarded as toroidal.
For other FRB sources in Section V.1.1, we adopt the parameters from Table 2 and $v_{\rm{w}}$ of $\sim 10^{7}-10^{8}$ cm/s to calculate $\mathrm{B}_{\rm ms}(r)$, $R_{\rm A}$ and the minimum $D_{\rm{\perp}}$ for each case (also see Table 3). Using a comparison of $R_{\rm A}$ of these sources to $D_{\rm{\perp,min}}$, we find that the FRB signals travel through the disk at locations well beyond $R_{\rm A}$ and in the toroidal magnetic field-dominated regime for each source. Therefore, in this work only the toroidal magnetic field of Equation (4) is considered.
The disk may also has a dipolar-like magnetic field of its own according to some researches (Balbus and Hawley, 1991; Dobler et al., 1999). However, given that the dipole magnetic field strength $\rm{B}_{\rm{d}}$ decays as $r^{-3}$, and the toroidal magnetic field strength $\mathrm{B}_{\mathrm{t}}\propto r^{-1}$, $\rm{B}_{\rm{d}}$ can be neglected compared to $\rm{B}_{\rm{t}}$ at the distance of $r\sim 10-10^{2}\,R_{\odot}$ (e.g., in the case of FRB 20201124A, using that $\rm{B}_{0}=10\,G$, $\rm{B}_{\rm{t}}$ is about 1 G at $D_{\rm{\perp,min}}=48\,R_{\odot}$, while $\rm{B}_{\rm{d}}$ is only about 0.01 G at the same distance).
V.2 Conclusions